OpenStudy (anonymous):

simplify 8P3. 42 336 40432 56

OpenStudy (warriorz13):

???

OpenStudy (warriorz13):

What does 8P3 mean?

OpenStudy (anonymous):

the numbers are smaller than the letter

OpenStudy (anonymous):

Warriorz isnt it lunch time?

OpenStudy (warriorz13):

por que?

OpenStudy (warriorz13):

All right @Alex_berry be back in 30 mins

OpenStudy (anonymous):

8p3 is (8!/(8-5)!) That simplifies to 8!/3! 8! is 8*7*6*5*4*3*2*1 = 40320 3! is 6. So the answer would be 40320/6 = 6720

OpenStudy (anonymous):

thats not right @sameerchandra

OpenStudy (anonymous):

because its not one of my answers

OpenStudy (anonymous):

@Alex_berry It should be. Maybe your answers are wrong. Once check if it is 8P3 or 8C3

OpenStudy (anonymous):

its 8P3

OpenStudy (anonymous):

Oh pellet I am sorry. It is 8!/5!. So the answer would be 336

OpenStudy (anonymous):

Am I right @Alex_berry

OpenStudy (anonymous):

i think so

OpenStudy (paki):

8P3= 336 @Alex_berry

OpenStudy (anonymous):

8p3 is (8!/(8-3)!) That simplifies to 8!/5! 8! is 8*7*6*5*4*3*2*1 = 40320 5! is 120. So answer is 336

OpenStudy (anonymous):

thanks

OpenStudy (paki):

pleasure